# Dense Feature Interaction Network for Image Inpainting Localization

### v1

@@ IV-C Qualitative Comparisons
In Fig. 7, we demonstrate the qualitative results on several image inpainting datasets. Our method achieves outstanding accuracy in localizing image inpainting across diverse manipulation regions and models, significantly surpassing the baseline methods. The segmentation results produced by DeFI-Net are closest to the ground truth, with rare occurrences of false positives. Specifically, DeFI-Net exhibits excellent performance when dealing with random regions (first and third columns) and specific semantic objects (second, fourth, fifth, and sixth columns) for inpainting. In contrast, MVSS-Net, IID-Net, and PS-Net give unsatisfactory results, with many mis-segmented regions and even overlooking the correct targets. PSCC-Net achieves decent results on some datasets, but it is prone to misjudgment in some challenging cases (such as slender, small targets) due to the influence of semantic information. In response to this issue, DeFI-Net obtains precise predictions thanks to our proposed dense feature pyramid learning strategy. This strategy delves into the potential relationships among multi-level features, better learning the inconsistencies in image details, structure, and semantics, enabling the model to response inpainting at different scales better. Moreover, the reverse edge attention enhancement effectively enhances edge features and suppresses the impact of semantic noise on inpainting representation, resulting in clearer boundary localization. Overall, DeFI-Net achieves satisfactory results.
@@ IV-D Computational Complexity Comparisons
To comprehensively evaluate the proposed method, we present comparisons of time consumption and computational complexity in Table V. All the compared methods are evaluated on the same device (GeForce RTX 3090 GPUs) with $256\times 256$ images. We can see that the proposed DeFI-Net introduces little computation overhead, and achieves comparable performance to existing models in terms of inference time.
TABLE V: Comparison of computational complexity and time consumption. The run time is averaged over 100 images.
@@ IV-E Visualization Results
To further present the discriminative and complementary features learned from various levels by DeFI-Net, we provide the visualization results of the feature maps at the low, middle, and high levels, as shown in Fig. 9. We can observe that the feature maps progressively capture the required discriminative information from low-level to high-level, including details, local structures, and global abstract semantics. Low-level features capture discriminative patterns around edges, middle-level features focus on texture inconsistencies in the filled areas, while high-level features identify changes in an image’s semantic content, such as localizing the entire objects. By effectively integrating these features, our model achieves precise localization across various inpainting regions and generation models.
@@ IV-F Ablation Studies
Essentially, DeFI-Net mainly benefits from three modules: DFPL, REAE, and SA-FF. Dense Feature Pyramid Learning (DFPL) highlights the specificity of features at different stages in the neural network, interacts quickly with neighboring layers in a densely connected manner, and improves the representation capability of features from other perspectives. Reverse Edge Attention Enhancement (REAE) enables the model to capture subtle differences at the boundaries by utilizing structural information provided by mid-level features as a priori guide. Spatial Adaptive Feature Fusion (SA-FF) adaptively learns how high-level features are fused with low and mid-layer features. To evaluate the effectiveness of DFPL, REAE, and SA-FF, we removed them separately from DeFI-Net and evaluated the inpainting detection performance on the IID-10K dataset with the ablation setup shown in Table VI. In addition, we performed ablation analysis for each module separately to demonstrate the effectiveness of their design.
Effect of DFPL. Previous works only use top-down [22] or bottom-up [57] for information interaction among features of different stages. While alleviating the efficiency problem of top-level or shallow-level information propagation in the opposite direction, it neglects the specificity of multilevel features. On the other hand, the densely connected method can fully utilize the adjacent and cross-layer information to enhance the context-awareness of features at each level. As can be observed from Table VII, compared with Var#1 (FPN [22]) and Var#2 (PANet [57]), DPFL significantly enhances performance and effectively promotes the spreading and utilization of multi-level features. An interesting phenomenon can be found from these results, where the performance of DPFL with concatenation fusion slightly outperforms that with addition fusion. The addition method may submerge shallow information, whereas dimension reduction with a $1\times 1$ convolution after concatenation can be viewed as a process of feature combination selection, which aids in the interaction of shallow features.
@@
TABLE VI: $F1$ and $AUC$ (%) of different combinations for our method on IID-10K dataset.
@@
TABLE VII: $F1$ and $AUC$ (%) of different variants for multi-level feature interaction on IID-10K dataset.
@@
Effect of REAE. Next, we evaluated the impact of the REAE module and presented the results in Table VIII. Comparisons with ordinary convolutional structure Var#1 and traditional edge extraction operator Var#2 verify the superiority of a priori learning, significantly improving the network’s performance. The selection of guidance features greatly affects the performance of the REAE. The results from Var#4 (namely our DeFI-Net model) demonstrate that using mid-level features as guidance yields the best performance. In fact, after DFPL, mid-level features fully integrate both shallow and deep information, and boundary learning leveraging REAE with strong contextual priors effectively suppresses noisy information.
TABLE VIII: $F1$ and $AUC$ (%) of different variants for edge supervision on IID-10K dataset.
@@
Effect of SA-FF. The fusion of high-level semantics and mid-low-level details is an effective method for enhancing localization accuracy. The SA-FF module dynamically allocates weights to different spatial positions simply. Table IX demonstrates the effectiveness of the SA-FF, where we compare different fusion methods, namely Var#1 (add fusion), Var#2 (concat fusion) and Var#3 (Dual attention [53]). The model can learn more universal fusion weights by adapting different attributes of specific inputs. Furthermore, we analyze the impact of different features used for learning weights. As indicated by Var#4 (mid-low-level) and Var#5 (high-level, namely our DeFI-Net model), the high-level features are more suitable for learning weights, probably due to their aggregation of more global information, which aids in understanding the broader context.
TABLE IX: $F1$ and $AUC$ (%) of different variants for feature fusion on IID-10K dataset.
@@
Effect of different loss terms. Furthermore, we thoroughly explored the effects of various loss terms. We have evaluated several parameter combinations for $\lambda_{r}$ and $\lambda_{e}$, corresponding to the coefficients of $L_{region}$ and $L_{edge}$, respectively. The results are shown in Table X. When $\lambda_{r}$ is greater than $\lambda_{e}$, the model’s performance improves. We ultimately select ($\lambda_{r}$=5, $\lambda_{e}$=1) as the default parameters for the loss function.
TABLE X: Ablation study with different loss weights of the loss function on IID-10K dataset.
@@ IV-G Failure Cases
Fig. 10 illustrates cases where our DeFI-Net does not yield accurate localization results, notably in scenarios featuring multiple tiny regions (the first two columns), images with in dark lighting conditions (the third column), and full-size synthesis (the last column). It is worth noting that even under these circumstances, our DeFI-Net still demonstrates highly competitive performance compared to existing methods.
@@ V Conclusion
In this work, we describe a new dense feature interaction network for image inpainting localization. Unlike existing methods, our model can effectively leverage the diverse and complementary information extracted from features at various stages within the backbone architecture. We employ a dense feature pyramid structure to capture multi-level and interactive traces left by inpainting. To mitigate the impact of semantic noise on the learning of artifact representation restoration, we propose a prior-guided edge enhancement strategy. Additionally, we present an effective spatially adaptive fusion technique for combining complementary representations, facilitating the adaptive fusion of mid-low-level features containing edge and shape information with high-level semantic features to achieve precise localization. Extensive evaluation demonstrates improved performance of our method. Our approach has limitations when handling large inpainting regions, particularly in fully manipulated images. In future works, we would like to investigate the generalization ability of more recent generative models. Additionally, it is important to investigate using adversarial training strategies to fortify our models against sophisticated post-processing or potential attacks.

### v2

@@ IV-C Qualitative Comparisons
In Fig. 7, we demonstrate the qualitative results on several image inpainting datasets. Our method achieves outstanding accuracy in localizing image inpainting across diverse manipulation regions and models, significantly surpassing the baseline methods. The segmentation results produced by DeFI-Net are closest to the ground truth, with rare occurrences of false positives. Specifically, DeFI-Net exhibits excellent performance in handling random regions (first and third rows) and specific semantic objects (second, fourth, fifth, and sixth rows) for inpainting. In contrast, MVSS-Net, IID-Net, and PS-Net produce unsatisfactory results, with many missegmented regions and even overlooking the correct targets. PSCC-Net achieves decent results on some datasets, but it is prone to misjudgment in some challenging cases (such as slender, small targets) due to the influence of semantic information. In response to this issue, DeFI-Net obtains precise predictions thanks to our proposed dense feature pyramid learning strategy. This strategy explores the potential relationships among multi-level features, better learning the inconsistencies in image details, structure, and semantics, enabling the model to better handle inpainting at different scales. Moreover, the reverse edge attention enhancement effectively enhances edge features and suppresses the impact of semantic noise on inpainting representation, resulting in clearer boundary localization. Overall, DeFI-Net achieves satisfactory results.
@@ IV-D Computational Complexity Comparisons
To comprehensively evaluate the proposed method, we present comparisons of time consumption and computational complexity in Table VI. All the compared methods are evaluated on the same device (GeForce RTX 3090 GPUs) with $256\times 256$ images. We can see that the proposed DeFI-Net introduces little computation overhead, and achieves comparable performance to existing models in terms of inference time.
TABLE VI: Comparison of computational complexity and time consumption. The run time is averaged over 100 images.
@@ IV-E Visualization Results
To further present the discriminative and complementary features learned from various levels by DeFI-Net, we provide the visualization results of the feature maps at the low, middle, and high levels. The intermediate features to be observed were iterated along the channel dimension and converted into PIL images using the torchvision.transforms.ToPILImage(). The results are shown in Fig. 9. We can observe that the feature maps progressively capture the required discriminative information from low-level to high-level, including details, local structures, and global abstract semantics. Low-level features capture discriminative patterns around edges, middle-level features focus on texture inconsistencies in the filled areas, while high-level features identify changes in an image’s semantic content, such as localizing the entire objects. By effectively integrating these features, our model achieves precise localization across various inpainting regions and generation models.
@@ IV-F Ablation Studies
Essentially, DeFI-Net mainly benefits from three modules: DFPL, REAE, and SA-FF. Dense Feature Pyramid Learning (DFPL) highlights the specificity of features at different stages in the neural network, interacts quickly with neighboring layers in a densely connected manner, and improves the representation capability of features from other perspectives. Reverse Edge Attention Enhancement (REAE) enables the model to capture subtle differences at the boundaries by utilizing structural information provided by mid-level features as a priori guide. Spatial Adaptive Feature Fusion (SA-FF) adaptively learns how high-level features are fused with low and mid-level features. To evaluate the effectiveness of DFPL, REAE, and SA-FF, we removed them separately from DeFI-Net and evaluated the inpainting detection performance on the IID-10K dataset with the ablation setup shown in Table VII. In addition, we performed ablation analysis for each module separately to demonstrate the effectiveness of their design.
Effect of DFPL. Previous works typically use either top-down [23] or bottom-up [62] for information interaction among features of different stages. While these methods alleviate the inefficiency of propagating top-level or bottom-level information in the opposite direction, they neglect the specificity of multilevel features. On the other hand, the densely connected method can fully utilize the adjacent and cross-layer information to enhance the context-awareness of features at each level. As can be observed from Table VIII, compared with Var#1 (FPN [23]) and Var#2 (PANet [62]), DFPL significantly enhances performance and effectively promotes the spreading and utilization of multi-level features. An interesting phenomenon can be observed: the performance of DFPL with concatenation fusion slightly outperforms that with addition fusion. This difference may arise because the addition method tends to suppress detailed information, whereas dimension reduction with a $1\times 1$ convolution after concatenation serves as a feature selection process, aiding in the interaction of low-level features.
@@
TABLE VII: $F1$ and $AUC$ (%) of different combinations for our method on IID-10K dataset.
@@
TABLE VIII: $F1$ and $AUC$ (%) of different variants for multi-level feature interaction on IID-10K dataset.
@@
Effect of REAE. Next, we evaluated the impact of the REAE module and presented the results in Table IX. Comparisons with ordinary convolutional structure Var#1 and traditional edge extraction operator Var#2 verify the superiority of a priori learning, significantly improving the network’s performance. The selection of guidance features greatly affects the performance of the REAE. Results from Var#4 (namely our DeFI-Net model) demonstrate that using mid-level features as guidance yields the best performance. After DFPL, mid-level features fully integrate detailed and abstract information. REAE, utilizing boundary learning with strong contextual priors, effectively suppresses noisy information.
TABLE IX: $F1$ and $AUC$ (%) of different variants for edge supervision on IID-10K dataset.
@@
Effect of SA-FF. The fusion of high-level semantics and mid-low-level details is an effective method for enhancing localization accuracy. The SA-FF module dynamically allocates weights to different spatial positions simply. Table X demonstrates the effectiveness of the SA-FF by comparing various fusion methods: Var#1 (add fusion), Var#2 (concat fusion) and Var#3 (Dual attention [58]). The model can learn more universal fusion weights by adapting different attributes of specific inputs. Furthermore, we analyze the impact of different features used for learning weights. As indicated by Var#4 (mid-low-level) and Var#5 (high-level, i.e., our DeFI-Net model), high-level features are more suitable for learning weights, probably due to their aggregation of more global information, which aids in understanding the broader context.
TABLE X: $F1$ and $AUC$ (%) of different variants for feature fusion on IID-10K dataset.
@@
Effect of different loss terms. Furthermore, we thoroughly explored the effects of various loss terms. We have evaluated several parameter combinations for $\lambda_{r}$ and $\lambda_{e}$, corresponding to the coefficients of $L_{region}$ and $L_{edge}$, respectively. The results are shown in Table XI. When $\lambda_{r}$ is greater than $\lambda_{e}$, the model’s performance improves. We ultimately select ($\lambda_{r}$=5, $\lambda_{e}$=1) as the default parameters for the loss function.
TABLE XI: Ablation study with different loss weights of the loss function on IID-10K dataset.
@@ IV-G Failure Cases
Fig. 10 illustrates cases where our DeFI-Net does not yield accurate localization, particularly in scenarios with multiple tiny regions (the first two columns), dark lighting conditions (the third column), and full-size synthesis (the last column). Notably, even under these challenging conditions, DeFI-Net still demonstrates highly competitive performance compared to existing methods.
@@ V Conclusion
In this work, we describe a new dense feature interaction network for image inpainting localization. Unlike existing methods, our model effectively leverages the diverse and complementary information extracted from features across various stages of the backbone architecture. We employ a dense feature pyramid structure to capture multi-level and interactive traces left by inpainting. To mitigate the impact of semantic noise on artifact representation learning, we propose a priori-guided edge enhancement strategy. Additionally, we present a spatially adaptive fusion technique to combine complementary representations, facilitating the adaptive fusion of mid-low-level features containing edge and shape information with high-level semantic features for precise localization. Extensive evaluations demonstrate the superior performance of our method.
Our approach has limitations when handling large inpainting regions, particularly in fully manipulated images. To address these challenges, our future research will focus on three key directions. First, we propose incorporating auxiliary tasks such as inpainting area classification, method recognition, and semantic anomaly detection to emphasize the distinctive global features between small-area inpainting, large-area inpainting, fully synthetic, and authentic images, thereby enhancing model generalization. Second, as noise pattern features are limited by susceptibility to post-processing, future work should establish a more robust noise representation through specialized neural networks and large annotated datasets to improve model adaptability. Finally, while pretrained vision models capture rich visual features and hold potential for transfer learning in inpainting detection, further research is needed to optimize architectures and training strategies for practical applications.